OpenStudy (anonymous):

4-9). Solve the following system of linear equations by elimination. x+y=10 -x+y=4 what is x? what is y?

OpenStudy (anonymous):

2y = 14 y=7 x = 10-7 x=3

OpenStudy (anonymous):

x+y=10 -x+y=4 cancel the -+(x) 2y=14 divide both side by 2. y=7 substitue the value of y to the first eq. x+7=10 x=10-7 x=3 sol.{x=3,y=7}

OpenStudy (radar):

Good answer awarded to rawxX45 for detailed and lucid answer.

OpenStudy (anonymous):

x+y=10 -x+y=4 solve for x x=10-y substitute this to the second -(10-y)+y=4 2y=14 y=7 x=10-7=3 Ans x=3, y=7